Circular No. 6601

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SAX J1750.8−2900

J. Marti, I. F. Mirabel and S. Chaty, Service d’Astrophysique, Centre d’Etudes de Saclay; and L. F. Rodriguez, Instituto de Astronomía, Universidad Nacional Autonoma de México, report: “We tentatively identify the radio counterpart of the galactic center x-ray transient SAX J1750.8−2900 with a variable VLA radio source, NVSS J175050−290311, the second of the three candidates listed on IAUC 6597. The 20-cm flux density of this object (which has positional uncertainty 10′′) was reported in the NRAO/VLA Sky Survey (NVSS) database to be 3.5 mJy (a measurement in 1993), whereas Helfand et al. (1992, Ap. J. Suppl. 80, 211) give a 20-cm flux density of 34 mJy (a measurement in 1989) for what is evidently the same radio source, 0.491−1.042, located at α = 17^{h}50^{m}51^{s}.47, δ = −29^{o}03′13′′.6 (equinox 2000.0, uncertainty 1′′).”

NOVALIKE VARIABLE IN SAGITTARIUS F. Kerber and H. Gratl, University of Innsbruck; and M. Roth, Las

Campanas Observatory, report: “We have found dramatic changes in the spectrum of the novalike object reported by Y. Sakurai in Feb. 1996 (cf. IAUC 6322, 6323) at α = 17^{h}52^{m}32^{s}.69, δ = −17^{o}41′07′′.7 (equinox 2000.0), compared to our observations from one year ago. The object was suspected to be undergoing a late He-flash (IAUC 6325, 6328), and our group has been monitoring the object’s evolution since discovery. Spectra obtained on 1997 Mar. 23.33 UT at the 2.5-m du Pont telescope show very prominent lines in the C_{2} Swan bands (bandheads 473, 516 and 558 nm) not seen before. In addition, the H Balmer lines seem to have weakened considerably compared to a year ago. The object was expected to change quickly in astronomical terms, but the observed change is certainly much faster and much more massive than expected.”

COMET 81P/WILD 2 Further to the report concerning comet C/1995 O1 on IAUC 6600,

J. Sarmecanic, D. J. Osip, M. Fomenkova and B. Jones report that midinfrared imaging of comet 81P with the 1.5-m reflector at Mt. Lemmon during Mar. 5.27–5.33 UT yields the following photometric magnitudes (10′′ circular aperture, uncertainties 0.2 mag): [8.7 μm] = 4.3, [10.3 μm] = 3.2, [11.7 μm] = 2.4 and [12.5 μm] = 2.1. The resulting color temperature is 200 ± 20 K, consistent with the equilibrium blackbody temperature of 213 K at r = 1.7 AU. There is no evidence for excess silicate emission.

1997 March 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden